OpenStudy (luigi0210):

Differentiate:

OpenStudy (luigi0210):

\[\LARGE y= (3x+1)^2\]

OpenStudy (anonymous):

2(3x+1)X3 =18x+6

OpenStudy (anonymous):

(dy/dx)=18x+6

OpenStudy (luigi0210):

Simplified, yes. Thank you :)

OpenStudy (owlcoffee):

Remember the derivative of a composite function: if we have a function the derivative is : \[y=g(x)^{m}\] \[y'=m.g(x)^{m-1}g'(x)\] So looking at the problem, it has pretty much the same structure: \[y=(3x+1)^{2}\] So following the derivative of a composite function (wich I didin't prove, just stated), the derivative should be, first derivating the function as if it was a single variable, and then multiplying by the derivative of the enclosed function: \[y'=2(3x+1)(3)\] applying distributive and multiplying we get: \[y'=(6x+2)(3)\] \[y'=18x+6\]

OpenStudy (tkhunny):

@harpreetsk Please never use "X" to indicate multiplication. VERY confusing. Usually "*" is clear enough. Oftentimes, simple juxtaposition is sufficient. Glad you're here on the job.

OpenStudy (anonymous):

thank for suggestion i will keep it in mind :)